40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 26 16:10:28 2009, Peter King, Configuration, PSL, PSL reference cavity temperature box modifications DSC_0731.jpg
    Reply  Wed Aug 26 16:35:03 2009, Alberto, Configuration, PSL, PSL reference cavity temperature box modifications 
    Reply  Sun May 8 22:59:40 2011, rana, Update, PSL, PSL reference cavity temperature box modifications 
       Reply  Mon May 9 20:30:20 2011, rana, Update, PSL, PSL reference cavity temperature box modifications Untitled.png
          Reply  Tue May 10 20:57:05 2011, rana, Update, PSL, PSL reference cavity temperature box modifications : after 24 hours still OK Untitled.png
Message ID: 1950     Entry time: Wed Aug 26 16:10:28 2009     Reply to this: 1953   4662
Author: Peter King 
Type: Configuration 
Category: PSL 
Subject: PSL reference cavity temperature box modifications 

The 40m Lab reference cavity temperature box S/N BDL3002 was modified as per DCN D010238-00-C.

These were:

    R1, R2, R5, R6 was 10k now are 25.5k metal film

    R11, R14 was 10k now are 24.9k metal film

    R10, R15 was 10k now are 127k thick film - no metal film resistors available

    R22 was 2.00k now is 2.21k

    R27 was 10k now is 33.2k

    U5, the LM-336/2.5 was removed

    An LT1021-7, 7 V voltage reference was added.  Pin 2 to +15V, pin 4 to ground, pin 6 to U6 pin 3.

    Added an 8.87k metal film resistor between U6 pin 1 and U4 pin 6.

    Added an 8.87k metal film resistor between U6 pin 1 and U4 pin 15.

    The 10k resistor between J8 pin 1 and ground was already added in a previous modification.

In addition R3, R4, R7, R8, R12 and R13 were swapped out for metal film resistors of the same value

(1.00k).

    The jumper connection to the VME setpoint was removed, as per Rana's verbal instructions.

This disables the ability to set the reference cavity vacuum chamber temperature by computer.

 DSC_0731.jpg

 

 

ELOG V3.1.3-